### v1

# Tree Attention: Topology-aware Decoding for Long-Context Attention on GPU clusters

Abstract
Self-attention is the core mathematical operation of modern transformer architectures and is also a significant computational bottleneck due to its quadratic complexity in the sequence length. In this work, we derive the scalar energy function whose gradient computes the self-attention block, thus elucidating the theoretical underpinnings of self-attention, providing a Bayesian interpretation of the operation and linking it closely with energy-based models such as Hopfield Networks. Moreover, due to this formulation, we discover that we can use efficient and optimized automatic-differentiation techniques to derive a highly efficient Tree Attention algorithm to compute the gradient of the energy and hence self-attention. Our formulation reveals that the reduction across the sequence axis can be efficiently computed in parallel through a tree reduction. Our algorithm, for parallelizing attention computation across multiple GPUs, enables cross-device decoding to be performed asymptotically faster (up to $8\times$ faster) than alternative approaches such as Ring Attention, while also requiring significantly less communication volume and incurring $2\times$ less peak memory. Our code is publicly available here: https://github.com/Zyphra/tree_attention.
1.  Introduction
*Denotes Equal Contribution
Correspondence to: vasu@zyphra.com, jonathan@zyphra.com
The self-attention operation is the core computational building block of the transformer architecture Bahdanau et al. (2014); Vaswani et al. (2017), which has become an ubiquitous and highly effective workhorse architecture currently applied at scale to language Brown et al. (2020); Kaplan et al. (2020); Hoffmann et al. (2022); Team et al. (2023); Achiam et al. (2023), vision Dosovitskiy et al. (2020), audio Betker (2023), and decision-making Chen et al. (2021); Reed et al. (2022). Nonetheless, the quadratic time complexity of self-attention means that significant resources are required to train and generate from Transformer-based Large Language Models (LLMs), especially for models with large context lengths.
During inference, the attention block largely determine the computational and memory requirements, which become more demanding as the input sequence length increases. Although LLMs generate one token at a time, the entire sequence of past tokens must still be stored in memory and used to compute attention scores during generation. Since attention performs a similarity matching of every token representation with every other, it incurs quadratic computational complexity in terms of flops. Moreover, it requires the entire sequence to be stored in memory during autoregressive generation.
While there has been recent advances in extending LLMs to extremely long contexts (up to 1M tokens)Chen et al. (2023); kai (2023); Peng et al. (2023), inferencing at such contexts lengths at scale still remains extremely expensive. This expense is unfortunate since long context models allow for qualitatively new capabilities such as extremely large scale in-context learning entire small datasets held in the prompt Reid et al. (2024); Lee et al. (2024); Bertsch et al. (2024), natural direct access to multimodal continuous data which is currently intractable for current transformers and which therefore must currently undergo a lossy tokenization scheme Reid et al. (2024); Team (2024), and potentially enables removing tokenization altogether and operating directly at the byte level Xue et al. (2022); Wu et al. (2024).
To speed up inference and alleviate memory requirements, recent works have attempted to alter the attention mechanism itself, either by linearizing it Katharopoulos et al. (2020), or approximating it by a kernel map Choromanski et al. (2020); Peng et al. (2021); Arora et al. (2024), which reduces the complexity to linear at the cost of reduced expressiveness. Others have invented alternative sequence mixing architectures such as state-space models which are designed to be efficiently computable in linear time and constant memory Gu and Dao (2023); Dao and Gu (2024); Katsch (2023); Sun et al. (2023); Glorioso et al. (2024). Alternately, other approaches utilize efficient algorithms to reduce the computational burden of attention while keeping the core computation the same. These include memory efficient attention Rabe and Staats (2021), Flash Attention Dao et al. (2022) and Flash Decoding fla (2024), which provide a set of IO-aware kernels to map the attention operation to the GPU hardware resources in an extremely efficient way, significantly reducing the memory overhead required. Further works, Character AI (2024); Kang et al. (2024); Liu et al. (2024); Nawrot et al. (2024) explore compressing or otherwise reducing the KV cache required in generation. Finally, Ring Attention Liu et al. (2023) proposes a way to parallelize the attention computation across the sequence axis between GPUs, thus enabling significantly longer contexts than can be served on a single GPU. This is the regime of primary interest of this paper. By leveraging the exact energy function for the self-attention block, we develop a method to speed up inference for long context use-cases when keys and values are sharded across multiple GPUs along the sequence axis.
Our proposed algorithm for computing attention via the gradient of the energy function is built on top of an efficient parallel computation and tree reduction communication strategy. In particular, this formulation lets us devise an asymptotically faster algorithm for performing decoding in which the number of communication steps scales logarithmically with the number of devices instead of linearly as alternatives such as Ring Attention Liu et al. (2023). Our topology aware approach illustrated in Fig. 1. significantly outperforms leading attention parallelization methods such as Ring Attention on multiple devices.
(a) Ring and Tree Attention Topologies. Due to the associative properties of the logsumexp and max operations of Tree Attention (Fig. LABEL:fig:tree), is possible to structure the reduction across the sequence as a tree, requiring asymptotically fewer communication steps than Ring Attention (Fig. LABEL:fig:ring) as well as less memory and communications volume.
In this work, we make three core contributions:
  • We provide a mathematical form for the energy function of self attention.
  • From this theory, we develop an algorithm for parallelizing the attention computation across devices, leveraging tree reduction topology.
  • Finally, we show empirically that our algorithm provides asymptotic speedups when decoding across many GPUs on long sequence lengths as well as large reductions in communication volume and peak memory usage.
2.  Related Works
Following the ubiquitous success of the transformer architecture, there has been significant effort to try to mathematically understand the nature and meaning of the attention operation and link it to existing and related literatures such as recurrent networks and Hopfield Networks. A number of recent works have attempted to study self-attention mathematically through the lens of energy functions. Ramsauer et al. (2020) pioneered this field by performing a similar but distinct analysis to relate self-attention with the modern Hopfield networks Krotov and Hopfield (2016), thus providing a novel and insightful interpretation of self-attention as performing hetero-associative memory lookups using a high-powered nonlinear similarity function Krotov (2021); Millidge et al. (2022). This work was later extended by Hoover et al. (2023) who derived a modified version of the transformer based off an energy function. Beyond this, other recent works attempt a Bayesian reformulation of attention by deriving a probabilistic generative model which matches the operations performed in a self-attention operation Singh and Buckley (2023).
However, while, it has long been known that the softmax operation can be derived as the gradient of the following scalar function:
$\partial_{z_{j}}\log\sum^{n}_{a=1}\exp(z_{a})=\frac{e^{z_{j}}}{\sum^{n}_{a=1}e^{z_{a}}}=\textrm{softmax}(z_{j}),$ (1)
known as the log-sum-exp, deriving an equivalent function for the self-attention block has not yet been performed.
In particular, consider the energy function proposed in Hoover et al. (2023) which was the logsumexp. Since the gradient in the update rule of that paper is taken with respect to the input to the block, the resulting function is a modified version of the self attention operation. Similarly, the update rule in Ramsauer et al. (2020) requires the tying of certain weights (K and V) in the attention operation. This restricts the Hopfield derivation to modelling autoassociative retrieval networks, while transformer attention is heteroassociative.
To advance beyond this, our energy function depends on a source tensor $\zeta$ (see section 4.), an auxiliary vector, and the gradient we take is with respect to this tensor. We will observe that this leads to the exact self attention operation when we take the gradient. Consequently, our formulation enables us to express self attention with no weight tying.
Another notable related work is Feng et al. (2024) where the authors also made similar observations as we do in section 5. about how the associative operations within the attention computation can be efficiently parallelized to motivate an attention based modified RNN architecture for sequence modeling.
While this energy function by itself is primarily a mathematical and theoretical curiosity, we demonstrate below that when combined with automatic differentiation, our formulation naturally leads to highly efficient parallel algorithms for computing attention and performing decoding, especially across multiple devices.
3.  Self-Attention
The self attention operation can be represented as a set of dot product similarity searches between queries and keys. These similarity scores are then reduced along the sequence axis and softmaxed, so that for a given query, there is a probability distribution of the similarities of each given key. We then take the expectation of the value vectors against this distribution. We denote the queries assigned to a sequence of length $N$ as $\{q_{a},a=1,\cdots N\}$, where each query is a vector of size $d$ that stands for hidden dimension: $q_{a}\in\mathbb{R}^{d}$ and similarly the keys and values $\{(k_{a},v_{a}),a=1,\cdots N\}$. Attention can be written as,
$z^{a}=\sum^{N}_{i=1}\text{softmax}(q_{a}\cdot k_{i}^{T})v_{i}$
The dot product, more explicitly reads:
$q_{a}\cdot k_{i}=\sum^{d}_{A=1}q_{a,A}k_{i,A},\forall A\in\{1,\cdots,d_{h}\}.$
The capital Latin indices ($A$) span the hidden dimension $d_{h}$. Naively computing attention in this way requires materializing the $qk$ matrix with quadratic computational and memory cost in the sequence length.
Memory efficient attention Rabe and Staats (2021) proposes an iterative way to compute the softmax similarities without ever having to materialize the full attention matrix. It performs the following operations, one query (or a chunk of queries) at a time:
$s^{(j)}_{i}$ $=\text{exp}(q_{j}\cdot k_{i})$ (2)
$n^{(j)}_{i}$ $=n^{(j)}_{i-1}+v_{i}s^{(j)}_{i}$ (3)
$d^{(j)}_{i}$ $=d^{(j)}_{i-1}+s^{(j)}_{i}$ (4)
Then, once the values $v$ and softmax denominator $d$ are computed we can divide to get the final softmaxed scores $z^{(j)}=\frac{n^{(j)}}{d^{(j)}}$ for every query index $j$. Since we compute attention in this iterative manner, the memory requirements are significantly reduced.
Flash attention Dao et al. (2022), utilizes a similar approach to reduce the memory and computational cost of attention but the algorithm is not adapted for multi-GPUs. Flash attention performs the iterative algorithm of Rabe and Staats (2021) in a blockwise manner, utilizing the block-parallel computational primitives available inside single GPU tensorcores. Additionally, it precisely sizes the blocks such that they can fit into the SRAM of the GPU for the entire attention computation, effectively performing kernel fusion and preventing many unnecessary IO operations.
4.  Self Attention as a gradient
Here, we will show how the self-attention operation can be written as the gradient of a certain scalar function. In particular, we will define a scalar function that depends on the keys, queries, values and additionally, on an auxiliary vector we will refer to as the source tensor $\zeta$. The source is the parameter with respect to which we compute the gradient of the scalar function to obtain the self attention operation. We need the source in order to write down the generating function of the moments of the distribution above. The source is the variable with respect to which we can Taylor expand the generating function and extract the moments as the coefficients of the monomials of $\zeta$ appearing in the Taylor series. Explicitly, we want to find a function $F(q,k,v,\zeta)$ such that:
$\sum^{N}_{a=1}\textrm{softmax}(q\cdot k_{a})v_{a}=\frac{\partial F}{\partial\zeta}\bigg|_{\zeta=0}.$ (5)
This terminology is inspired by work on energy-based models in machine learning Beal (2003); LeCun et al. (2006); Song and Kingma (2021). A summary of variables and indices is provided in appendix  11.
4.1.  The energy function
In what follows, we will show how the energy function we are after is given by the generating function for the cumulants of the distribution given by attention scores. Taking inspiration from statistical mechanics where an analogous cumulant generating function defines the Helmholtz Free energy Wikipedia (2024), we dub our cumulant generating function the energy function for self attention.
Let us focus on the case where there is a single query only. As noted above, we leverage the fact that the attention operation can be seen as the computation of the expectation value of the vectors $v$ in the distribution set by the attention scores $z$:
$z=\langle v\rangle=\sum^{N}_{a=1}P_{a}v_{a}=\frac{\sum^{N}_{a=1}e^{q\cdot k^{T}_{a}}v_{a}}{\sum^{N}_{i=1}e^{q\cdot k^{T}_{i}}}.$ (6)
We see that the probability density is given by:
$P_{a}=\frac{e^{q\cdot k^{T}_{a}}}{\sum^{N}_{i=1}e^{q\cdot k^{T}_{i}}},$ (7)
and satisfies:
$\sum^{N}_{a=1}P_{a}=1.$ (8)
Typically, the denominator or normalization factor is identified with the so called partition function:
$Z=\sum^{N}_{a=1}e^{q\cdot k^{T}_{a}}.$ (9)
We can now compute the first moment of the probability distribution given above by introducing a source, $\zeta\in\mathbb{R}^{d}$. In our case, with $\zeta,$ we can extend the partition function to the function:
$Z(\zeta)=\sum^{N}_{a=1}e^{q\cdot k^{T}_{a}+\zeta\cdot v^{T}_{a}}.$ (10)
Now, we can compute any moment of the distribution by as the $n$th Taylor coefficient of $Z(\zeta)$ $\forall A_{1},A_{2},\cdots\in\{1,\cdots,d_{h}\}$ :
$\langle v_{A_{1}}\cdots v_{A_{n}}\rangle=\frac{1}{Z}\frac{\partial^{n}Z(\zeta)}{\partial\zeta_{A_{1}}\cdots\partial\zeta_{A_{n}}}\bigg|_{\zeta=0}.$ (11)
In other words, we can write $Z(\zeta)$ as:
$Z(\zeta)=Z\left(1+\langle v\rangle\zeta+\frac{1}{2!}\langle v_{A_{1}}v_{A_{2}}\rangle\zeta_{A_{1}}\zeta_{A_{2}}+\cdots\right)$ (12)
So we see that the first moment can be written as:
$\langle v\rangle=\frac{1}{Z}\frac{\partial Z}{\partial\zeta}\bigg|_{\zeta=0},$ (13)
which can be written as the gradient of the log of $Z(\zeta)$:
$\langle v\rangle=\frac{\partial}{\partial\zeta}\log Z(\zeta)\bigg|_{\zeta=0}.$ (14)
This quantity is what we call the generating function, or the free energy:
$F=\log\sum_{a}\exp\left(q\cdot k_{a}^{T}+\zeta\cdot v^{T}_{a}\right).$ (15)
To compute causal self attention, then we introduce $N$ sources $\zeta^{i}$ each $\in\mathbb{R}^{d}$ and take
$F_{tot}=\sum^{N}_{i=1}F_{i}=\sum^{N}_{i=1}\log\sum^{i}_{a=1}\exp(q_{i}\cdot k^{T}_{a}+\zeta_{i}\cdot v^{T}_{a}).$ (16)
The truncation of the inner sum up to index $i$ is due to causal masking.
[…]
$\frac{\partial F_{tot}}{\partial\zeta_{i,A}}\bigg|_{\zeta_{i}=0,\forall i}=\frac{\sum^{i}_{a=1}\exp(q^{i}\cdot k^{T}_{a})v_{a,A}}{\sum^{i}_{a=1}\exp(q^{i}\cdot k^{T}_{a})}.$ (17)
The generalization to the multi-head attention case is straightforward. There we have one key, query and value per head. For $n_{h}$ total heads, the generating function takes the form:
$F_{tot}=\sum^{N}_{i=1}\sum^{n_{h}}_{h=1}F^{i,h},$ (18)
where
$F_{i,h}=\log\sum^{i}_{a=1}\exp\left(q_{i,h}\cdot k_{h,a}^{T}+\zeta_{h,i}\cdot v^{T}_{h,a}\right).$ (19)
The output projection weight is included in the definition of $v_{j}$ here, meaning that
$v_{b,A}=x_{b,\bar{B}}(W_{O}W_{V})_{A,\bar{B}}$ (20)
where $W_{O}\in\mathbb{R}^{d_{h}}\times\mathbb{R}^{d_{emb}}$ denotes a head size slice of the output projection weight and $\bar{B}\in\{1,\cdots,d_{h}\}$ spans the intra-head indices. In the index notation above, the head indices are barred whereas the embedding space indices are unbarred. We will proceed further just focusing on the single head case as it makes the presentation more simple and the multi-head generalization is immediate. Note that we demonstrate that our energy function approach also can account for safe softmax in Appendix 10.
4.2.  Bayesian interpretation
The fact that it is possible to derive the self-attention operation as the minimization of an energy function implies that it is possible to provide a Bayesian gloss on self-attention by identifying a likelihood function and show that we can obtain the forward pass of the attention block from computing the maximum a posteriori estimate of this likelihood.
In particular, we propose the following for the log-likelihood function:
$\Gamma(\zeta,z)=\sum^{N}_{i=1}\sum^{d}_{A=1}\bigg(z_{i,A}\zeta_{i,A}-F(\zeta,x)\bigg).$ (21)
We denote by $x$ the input to the self attention block from which we obtain $q,k,v$ from multiplying it by the weights $W_{Q},W_{K},W_{V}$ respectively. Let us minimize the above with respect to $\zeta$ and $z$ simultaneously:
$\frac{\partial\Gamma}{\partial\zeta_{i,A}}=0,$ (22)
$\frac{\partial\Gamma}{\partial z_{i,A}}=0.$ (23)
These conditions written explicitly read
$\zeta_{i,A\,*}=0,\,\,\,\,z_{i,A\,*}=\frac{\partial F}{\partial\zeta_{i,A}}.$ (24)
Plugging the first condition into the second leads to the attention forward pass:
$z_{i\,*,A}=\frac{\sum^{i}_{a=1}e^{q_{i}\cdot k^{T}_{a}}v_{a,A}}{\sum^{i}_{b=1}e^{q_{i}\cdot k^{T}_{b}}}.$ (25)
In all, this means we can obtain the gradient w.r.t. $\zeta$ from MAP estimation of the following likelihood:
$z_{i\,*,A},\zeta_{i\,*,A}=\textrm{argmax}_{\zeta,z}e^{-\Gamma(\zeta,z)}.$ (26)
Moreover, such a procedure enables us to identify the energy based model associated with the self-attention function.
5.  Tree Attention
In this section we will show how the formulation of the attention operation as the gradient of an energy function suggests an efficient parallel strategy for computing it. The key insight is to leverage an efficient algorithm to compute the energy, and then differentiate it in order to obtain an efficient algorithm to compute attention.
5.1.  Computing the energy efficiently
Let us focus on the case of decoding with a KV cache in a causal language model where we have one query and $N$ keys and values. In this case, the energy function is:
$F_{dec}=\log\sum^{N}_{a=1}\exp(q\cdot k_{a}^{T}+\zeta\cdot v^{T}_{a})$ (27)
$\equiv\textrm{logsumexp}_{a}(\{q\cdot k^{T}_{a}+\zeta\cdot v^{T}_{a},a=1,\cdots,N\}).$ (28)
A crucial and important fact is that both $\textrm{logsumexp}_{a}$ and $\max_{a}$ are associative operations:
$\textrm{logsumexp}_{a}(\{T_{a},\textrm{logsumexp}_{a}(\{R_{a},S_{a}\})\})=$
$\textrm{logsumexp}_{a}(\{\textrm{logsumexp}_{a}(\{T_{a},R_{a}\}),S_{a}\}),$
$\max_{a}(\{\max_{a}(\{T_{a},R_{a}\}),S_{a}\})=$
$=\max_{a}(\{T_{a},\max_{a}(\{R_{a},S_{a}\})\}).$
We can prove that this associative property allows these reduction to be performed efficiently in parallel with logarithmic time complexity, provided we have adequately many parallel workers:
Theorem 1.
The time complexity of a reduction operation involving an associative function, such as $\textrm{logsumexp}_{a}$ or $\max_{a}$, over an array of size $N$ using $p$ parallel processors is $O\left(\frac{N}{p}+\log p\right)$. When the number of processors $p$ is equal to $N$, the time complexity is reduced to $O(\log N)$.
[…]
Algorithm 1 Single Query Energy Forward
[…]
2:Scatter a copy of $\mathbf{q},\mathbf{\zeta}$, and each $\mathbf{k}_{\hat{a}},\mathbf{v}_{\hat{a}}$ to each of the $p$ processors.
3:In parallel compute $r_{\hat{a}}=\mathbf{q}\cdot\mathbf{k}^{T}_{\hat{a}}+\mathbf{\zeta}\cdot\mathbf{v}^{T}_{\hat{a}}$
4:Compute $m=\textrm{Reduce}(\max,r)$ by doing a tree reduction.
5:Scatter $m$ to every device and update $r_{\hat{a}}\rightarrow r_{\hat{a}}-m$.
6:Compute $O=\textrm{Reduce}(\textrm{logsumexp},r)$ by doing a tree reduction.
7:Save $O,m$ for gradient w.r.t $\zeta$.
8:Return $O$
5.2.  Efficient parallel decoding
One of the core insights of automatic differentiation is that the gradient of a function $\nabla_{x}f(x)$ can be computed with the same time complexity as computing $f(x)$ Vieira (2016). The caveat however is that if the function has a deep computational graph, then the memory footprint of computing the gradient grows with that depth as backpropagation requires storing the values of the intermediate tensors. In our case, the computational graph involved in computing the energy is shallow and therefore the memory overhead is negligible. This means that if we can compute the energy efficiently, we obtain an efficient algorithm for computing its gradient (i.e. the self-attention operation) automatically.
In our case, we want to compute the gradient of the energy function with respect to $\zeta_{A}$ and then set it to zero. This can be done with automatic differentiation engines having set $\zeta$ to be a tensor of zeros from the very outset. We can however manually implement a gradient with respect to $\zeta$ pass of the above Algorithm  1 that doesn’t materialize $\zeta$ in Algorithm 2 below. Note in particular that when we set $\zeta_{A}=0,\,\,\forall A\in\{1,\cdots,d_{h}\}$ then $O$ involves only the logsumexp of the dot product between queries and keys.
Algorithm 2 Single Query Energy Gradient w.r.t $\zeta$
[…]
2:Scatter a copy of $\mathbf{q},m$ and $O$, and each $\mathbf{k}_{\hat{a}},\mathbf{v}_{\hat{a}}$ to each of the $p$ processors.
3:In parallel compute $r_{\hat{a}}=\mathbf{q}\cdot\mathbf{k}^{T}_{\hat{a}}-m$
4:Compute $R_{\hat{a}}=\exp(r_{\hat{a}}-O)\cdot v_{\hat{a}}$
5:Compute $z=\textrm{Reduce}(\textrm{sum},R)$
6:Return $z$
Notice here that by storing $O,m$ for backward, the onlt remaining reduction operation that needs to be performed is the one in line 5 of the above algorithm. This single reduction takes $O(N/p)$ time to compute the local sums on each device and $\log p$ time to communicate and combine partial results therefore we get the same asymptotic complexity as the logsumexp calculation.
In practice we implement the forward and gradient w.r.t. $\zeta$ in a single function which returns both the value and the gradient of the energy function. We can therefore put together Algorithms 1 and 2 into the following efficient parallel decoding algorithm 3:
Algorithm 3 Tree Decoding
[…]
2:Scatter $\mathbf{q}$ to each GPU.
3:On each GPU, use Flash Attention2 to compute $\textrm{o}=\frac{\sum_{\hat{a}}\exp(\mathbf{q}\cdot\mathbf{k}_{\hat{a}}^{T})\mathbf{v}_{\hat{a}}}{\sum_{\hat{b}}\exp(\mathbf{q}\cdot\mathbf{k}^{T}_{\hat{b}})}$ and $\textrm{lse}=\log\sum_{\hat{b}}\exp(\mathbf{q}\cdot\mathbf{k}^{T}_{\hat{b}})$.
4:Obtain the local numerator and denominator by then computing: $\textrm{n}=\textrm{o}*\exp(\textrm{lse}),\textrm{d}=\exp(\textrm{lse})$.
5:Recompute the global max $m=\textrm{{Allreduce}}(\textrm{max},\textrm{d})$.
6:Compute global numerator and denominator $\textrm{n}_{g}=\textrm{{Allreduce}}(\textrm{sum},\textrm{n}e^{-m}),\textrm{d}_{g}=\textrm{{Allreduce}}(\textrm{sum},\textrm{d}e^{-m})$.
7:Return result $z=\frac{\textrm{n}_{g}}{\textrm{d}_{g}}$.
This algorithm requires three Allreduce operations in total which means that the time complexity it requires is $O(3(N/p+\log p))$.
5.3.  Implementation and topology-awareness
Figure 2: NCCL Send/Recv between two H100 GPUs intra-node and inter-node. GPU clusters offer a two-tier topology whereby intra-node bandwidth is significantly higher than inter-node. Algorithms such as Tree Attention exploit this topology by reducing internode communication requirements, enabling better overlap of communication with computation.
(a) Execution time of of 16 head Tree Attention vs Ring Attention for different sizes of GPU cluster (from 1 to 16 H100 DGX nodes). Relative Execution Times are indexed to the Ring Attention times at a sequence length of 80k tokens.
While the theoretical analysis above indicates that we should see speedups when using tree-based reductions, this is not necessarily guaranteed in practice due to various potential overheads. However, importantly, beyond its asymptotic benefits, Tree Attention benefits from taking advantage of the two-level topology which is standard in modern GPU clusters.
Ring Attention is inherently not topology-aware, and only scales within a network of homogeneous bandwidth. However, this is in conflict with the two-level network topology of modern GPU clusters, which use high-bandwidth interconnects within nodes (NVLINK or PCIe) and comparatively lower-bandwidth interconnects across nodes (InfiniBand or Ethernet). The interconnects greatly differ in bandwidth and latency (see Figure 2). Therefore, Ring Attention is bottlenecked by the slowest interconnect, and cannot always overlap the attention computation with communication. We discuss this point further in 6.3. Tree Attention improves on Ring Attention by using network topology-aware communication patterns to increase overlap of computation and communication, and decrease this scalability bottleneck on communication from the distributed attention computation.
In practice, collective communication libraries like NCCL attempt to automatically detect what the right communication strategy is based on considerations such as data volume and network topology. In DGX clusters, for collective operations within a node, ring reduce is performed whereas a tree reduction is performed across nodes. We see that therefore using built-in collective operations such as Allreduce leads to a better performance when decoding from long contexts across multiple GPUs than enforcing the Ring Attention’s point to point communication pattern. We show how the following strategy outperforms Ring Attention when decoding from very long contexts across multiple GPUs.
In our empirical experiments , we use Flash Attention 2 Dao (2023) within each device, both for our algorithm and for Ring Attention [^1] A JAX based Ring Attention implementation that uses Flash Attention 2 can be found here: https://github.com/nshepperd/flash_attn_jax.. We provide a simple JAX implementation of our method in Appendix 8.. Note that our method mirrors Flash Decoding fla (2024) except in that case, the parallelization happens at the level of different streaming multiprocessors (SMs) within a GPU whereas we parallelize between different GPUs. All computations are performed in BF16.
6.  Results
Similar to Ring Attention, Tree Attention is an exact computation of Attention. Since training and evaluation metrics are the same as Attention, our experimental results are focused primarily on latency in section 6.1., peak memory usage in section 6.2. and communication volumes in section 6.3.. Since our algorithm computes numerically identical results as the forward pass of standard Attention, our performance results transfer seamlessly to Transformer architectures.
We performed all experiments on a DGX H100 cluster consisting of 16 nodes, each containing 8 H100 GPUs. All GPUs within the node are connected via an all-to-all NVLINK 4.0 (900GBps) topology. Nodes are connected to each other via 8 InfiniBand NDR interconnects per node (1 per GPU), each of which provides 400 Gbps (leading to an aggregate 3.2 Tbps node injection bandwidth).
6.1.  Latency
In terms of practical usefulness, our study of the energy function brought to light a previously unnoted parallelizability inside the attention computation – that of the reduction of the logsumexp across the sequence dimension, which can be implemented as a parallel Allreduce. As stated in Theorem 1, it becomes theoretically possible to implement attention, per query as an $N/p+\log(p)$ parallel operations rather than $N$, where the logarithmic term is proportional to the number of devices available for parallelization. When the attention is sharded across multiple devices, this asymptotic speedup creates a considerable speedup over alternative methods for decoding.
To empirically test the theoretical benefits of our Tree Attention method, we compute latency by measuring the time required to perform decoding for different sequence lengths and varying number of H100 nodes. We compare Tree Attention to our own Ring Attention execution times in Fig. 4(a). Both methods use FlashAttention-2 Dao (2023) for the individual-GPU attention computation. For our experiments, we benchmark on a standard attention block consisting of 16 heads of dimension 128 across different sequence lengths.
Our latency results shows how Tree Attention improves over Ring Attention as we increase the sequence length in Fig. LABEL:fig:relative_times and increase the number of GPUs in Fig. LABEL:fig:absolute_times. As the plots demonstrate, as we scale the sequence length or the number of GPUs, the gap between Tree Attention and Ring Attention execution time widens asymptotically. Remarkably, Tree Attention achieves close $\times 8$ speedups when we use 128 GPUs on a sequence length of 5.12M. We expect this trend to continue for larger sequence lengths.
6.2.  Memory cost
(b) Peak memory usage of a single attention block with Tree Attention vs Ring Attention when sharded between two RTX 4090s. Results taken using the JAX memory profiler on one GPU. The difference in peak memory scales with hidden size and sequence length.
[…]
$\textrm{Mem}_{\text{ring}}$ $=4btd+2bd$ (29)
$\textrm{Mem}_{\text{tree}}$ $=2btd+2bd+2bn_{h},$ (30)
where $d=d_{h}\times n_{h}$, for head size $d_{h}$ and $n_{h}$ number of heads, $b$ denotes the batch size and $t=N/p$. As such, so long as $2bn_{h}\leq 2btd$, which will almost always be the case in realistic scenarios, our method always has a lower peak memory cost compared to Ring Attention.
We empirically measure peak memory utilization for our approach and Ring Attention to show that indeed memory usage is significantly less for Tree Attention in Figure 4(b). As predicted by theory, scaling hidden size or sequence length scales Ring Attention peak memory usage about $2\times$ faster than Tree Attention. For example, doubling the hidden size from 2048 to 4096, doubles the gap in peak memory between two methods, going from 524MB to 1040MB.
6.3.  Communication volume
For Ring Attention’s P2P communication strategy, the total volume of data being communicated between devices (in units of number of tensor elements) per iteration is given by:
$V_{ring}=2btd\times p$ (31)
where $p$ is the number of devices. The first factor comes from counting the total number of communicated elements corresponding to $\{(\mathbf{k}_{\hat{a}},\mathbf{v}_{\hat{a}}),\hat{a}=1,\cdots,t\}$, i.e.
$\textrm{numel}\left(\{(\mathbf{k}_{\hat{a}},\mathbf{v}_{\hat{a}}),\hat{a}=1,\cdots,t\}\right)=2btd.$ (32)
The Allreduce strategy we use in Tree Decoding requires the following volume Anthony et al. (2024):
$V_{\texttt{Allreduce}}=2\times\frac{p-1}{p}\times\textrm{numel}.$ (33)
We communicate a shard of the numerator, denominator and max, requiring:
$\textrm{numel}\left(\textrm{n},\textrm{d},m\right)=bd+2bn_{h}.$ (34)
Note that we first perform on device the local reductions to obtain the local numerator and denominator on each device which consequently makes it so that $t$, i.e. the size of the local sequence chunk doesn’t appear in the above expression. We then obtain:
$V_{Tree}=2\frac{p-1}{p}\times\left(bd+2bn_{h}\right).$ (35)
Our theoretical analysis shows that per iteration our algorithm maintains a lower communication volume than Ring Attention. Note however that Ring Attention when performed in the training setting with many queries overlaps communication and computation so as to hide it’s communication costs. However, overlapping communication and computation in the decoding case is infeasible because of how fast the attention computation on a single GPU is relative to how long it takes to communicate the chunk of keys and values between two devices.
Concretely, let us take the example of decoding from a context of length 640000 split between $8$ GPUs within one node. Let us take a hidden size of 2048 and fix our data type to be bfloat16. Each device for decoding takes $O(10^{-5})$ seconds to perform the Flash Attention computation. The time it takes to move the keys and values of the corresponding size between adjacent GPUs as per Fig. 2 is roughly $O(10^{-3})$ seconds. The latency incurred between nodes is even greater and therefore overlapping isn’t feasible due to this disparity in timescales.
7.  Discussion and Conclusion
In this paper, we have derived the energy function for self-attention and demonstrated how the computation of the derivative of this function provides a novel and efficient method for computing attention in parallel. This advantage is especially apparent when performing decoding across multiple devices, in which case our Tree Attention enables us to substantially outperform Ring Attention with an asymptotically superior algorithm, with $\times 8$ speedups when we use 128 GPUs on a sequence length of 5.12M. We also see that the AllReduce operation that we use involves sending partially reduced objects, which greatly reduces the volume of communicated data as well as the peak memory requirement.
Our introduction of a unique energy function for self attention develops interesting connections between attention and other related models such as Hopfield networks and the general notion of associative memories. Given this energy function, it is possible to mathematically analyze its loss landscape and dynamics, as well as begin to understand how the attention operation could potentially be improved. Additionally, since the energy function can be framed as a variational free energy, it becomes possible to interpret the attention operation through a Bayesian lens and understand the generative model that is implicit in its operation. We hope that these mathematical derivations aid research in understanding the properties of transformer networks, as a deep understanding of the Hopfield energy function has done for Hopfield networks.
Theoretical asymptotic improvements should be demonstrable in the single-device case, but in practice, this advantage is not realized. SMs in H100 and older architectures communicate through shared global memory, not in a peer-to-peer fashion like individual GPUs in a multi-device setup, making the reduced communication negligible. However, recent experimental instructions in the H100 have enabled peer-to-peer SM communication  NVIDIA (2024), suggesting that clever use of these instructions could lead to speedups over Flash Attention on a single device.